「レセプションパーティへのお誘い」 (平井武)

Invitation to Reception Party on Thursday, the 21st February, Evening 18:00 – 20:30 prepared by Mr. and Mrs. HIRAI,

Dear Collleagues, Participants of the workshop held at RIMS, 18th – 22nd, February

We, Mr. and Mrs. HIRAI, will prepare a reception party for all the participants of this workshop, in the evening of Thursday, the 21st February, in the campus of Faculty of Integrated Human Studies.

We cordially invite you to this informal party, and we will be very happy if you could join us.

No formal speaches are assumed, and you can just enjoy freely foods, drinks and conversations.

The details are as follows.

Date and Time: Thursday, the 21st February, Evening 18:00 〜 20:30

Place: 4th floor, Seminar Room #1401(Room Tel: 753-7895) in the building No.1 at the left side of the main gate

1 号館 4 階セミナー室， 1401 号セミナー室 (1401 セミナー室の電話番号は 753-7895)

《レセプション会場が 2 階だとご案内した方もありましたが建物を間違えていました．そして， 4 階でした》

Caution: The doors of the entrance of this building will be closed at about 7 o’clock. After this time you cannot open the door from the outside (it can be opened from the inside).

In case you have some problems, please call the phone number indicated above.

表現論でなくても関係のありそうな方，院生の方もどうかお誘い合わせ，おいで下さい．

平成 13 年 2 月 13 日 (13/02/2002) 平井武，悦子 (Takeshi and Etsuko Hirai)